The density of a certain gas is 1.98gdm-3 at s.t.p. What is the molecular mas of the gas?
A. 44.0 g
B. 54.0 g
C. 26.0 g
D. 31.0 g
Show working
1
Expert's answer
2020-04-18T07:01:17-0400

D=1.98kgdm-3 at s.t.p

at s.t.p; T=273.15K, P=100000pa

"V=RT\/P"

V=(8.314*273.15)/100000=0.0225m3

mass,"m=DV"

m=1.98*0.0225 =0.044kg

molecular mass,m =A. 44g